### v1

@@ 3 Methodology
3.1 Data Collection
In this study, data are collected from a publicly accessible source known as the "Monkeypox Skin Lesion Dataset" R26. The dataset includes two classes: Monkeypox and Non-Monkeypox, which are used for performing binary classification. The Monkeypox class contains 1428 skin images of monkeypox lesions. On the other hand, Non-Monkeypox class contains 1764 skin images of Non-Monkeypox lesions. The detailed data description is shown in Table 2. Figure 3 illustrates a visual representation of the dataset’s image distribution, highlighting the proportion of monkeypox lesion images compared to non-monkeypox lesion images. This information is crucial for understanding the composition of the dataset and the relative representation of different lesion types, thus providing valuable insights into the data used for training and evaluation in our study. Figure 4 shows the Sample set of images of Monkeypox lesion and Non-Monkeypox lesion from the dataset.  [^1] https://www.kaggle.com/datasets/nafin59/monkeypox-skin-lesion-dataset
Table 2: Dataset Description.
@@ 3.1 Data Collection
Monkeypox In this dataset, the monkeypox class contains 1428 skin images of monkeypox lesions. Kaggle Datasets R26
Non-Monkeypox Non-Monkeypox class contains 1764 Skin images of chickenpox lesions, measles lesions. Kaggle Datasets R26
@@ 3.2 Data Pre-processing
This step is crucial for the deep learning model as it ensures that the input data is appropriately formatted for training the model, leading to enhanced model accuracy. In this step, collected data are passed through six pre-processing operations before it is fed into the model. These pre-processing operations include resizing images to uniform sizes, rescaling the pixel value, setting data_format, label encoding, in-place augmentation, and data new3. All pre-processing operations are described here.
3.2.1 Resizing Image
@@
Image resizing is the process of resampling or downsampling the image. Images in the dataset should have the same size as the CNN model. So, these stages are used to assure that all the images have the same size. This stage is also used to reduce the computational burden of training the model. This is accomplished with the aid of Python/TensorFlow.keras.preprocessing module. In this study, image dimensions are set to 224×224 dimension. The reason for selecting image size 224×224 is that this size allows better performance, more accurate training, and faster convergence of model R28.
3.2.2 Rescale:
@@
Rescaling is the process of adjusting or normalizing the image data to a specific range of 0 to 1. ImageDatagenerator class is used to transform the pixel values of images from the original range of 0 to 255 into the desired range of 0 to 1, which is more suitable for neural network models. Different image has different pixel value. If rescale operation is not applied, the model only encounters certain images which leads to overfitting. This operation ensures that each image is given equal importance and weight during the model’s training process. This operation can be accomplished by specifying the rescale argument which involves multiplying each pixel by a specific ratio to achieve the desired scaling. In this scenario, the ratio used for rescaling the pixel values is 1/255, which is approximately equal to 0.0039 R24.
3.2.3 Data_format:
@@
In this study, data_format is set to "channels_last" which means that the last dimensions of the input tensor are channels. In this "channels_last" format image is depicted as a 3D tensor that corresponds to (height, width, and channels). In deep learning frameworks like TensorFlow and Keras, this format is the default value in Keras. During training and evaluating the model, this format ensures compatibility, efficient computation, and data manipulation R29.
3.2.4 Label Encoding:
@@
Label encoding is a process of transforming categorical labels into numerical representations, enabling deep-learning predictive models to comprehend them effectively. In this study, the dict and zip method is used to create a mapping between the two classes ‘Monkeypox’ and ‘Non_Monkeypox’ and their corresponding numerical values 0 and 1. By using the dict method, we have created a dictionary in which the keys are the numerical values and the values are corresponding categorical labels. The zip method then makes a list of rows by combining the numerical values with the categorical labels. With the resulting dictionary, the categorical labels can be transformed into numerical values, allowing the model to utilize these numerical representations for making predictions. R27.
3.2.5 Data Augmentation (in-place):
@@
In this pre-processing operation, random image transformations performed by in-place data augmentation (on-the-fly) have been applied, which means that the dataset size does not increase. In this type of data augmentation, at the time of training, our network observes unique variations of our data during each epoch and captures different patterns and features with each iteration. To transform the image, ImageDatagenerator, a Keras image prepossessing library, is used. Image transformation operations available in the ImageDatagenerator function include brightness_range, zoom range, horizontal flip, etc. Table 3 depicts the augmentation technique used in this study. Figure 5 shows the in-place augmentation process R30. The process of applying in-place data augmentation is R30: Step-1: At first, augmentation parameters including zoom_range, brightness_range, fill_mode, and horizontal_flip are defined. Step-2: An instance of an ImageDataGenerator class, specifying the augmentation parameters is also created. Step-3: An original batch of images is passed to the ImageDataGenerator class. Step-4: Then, the batches of augmented data are generated and returned to the calling method. Step-5: After that, the model is trained on batches of images.
@@
Table 3: Augmentation technique used in this study.
@@
3.2.6 Data Visualization:
This stage requires sampling and reviewing the input data to observe the randomly transformed images and to verify that the transformations are being applied as expected. This stage is also used to confirm the readability of the input images. ImageDataGenerator library is used to generate a batch of randomly transformed images to visualize the transformed images. These transformed images are then visualized using the visualization library Matplotlib via 2D representation with the new picture dimensions R27. Figure 6 shows the visualization of the transformed image sample of our dataset.
3.3 Splitting Dataset
@@ 3.4 Model Building
Table 4: Summary of models used in this study.
@@
3.4.1 Architecture of our Proposed MpoxSLDNet model :
@@
Figure 12 shows the graphical representation of fully connected layer. Then the last layer is 2 units dense layer with a sigmoid activation function producing a binary classification output. Table 4 shows the summary of hyperparameters used in the proposed MpoxSLDNet model.
Table 5: Hyper parameters for the proposed MpoxSLDNet model.
@@ 4 Result Analysis
4.1 Assessing the MpoxSLDNet model’s performance through multiple runs
Table 6 shows the training and validation accuracy obtained from running the proposed model three times for 20 epochs. The purpose of multiple runs is to mitigate the effects of randomness during the training phase and provide a more reliable evaluation of how well the model performs. Based on the Table 6, the MpoxSLDNet model attained an average training accuracy of 96.89% and an average validation accuracy of 93.08%. While the performance of the model is consistent across the three runs, there is some variation in the accuracy values. This highlights the importance of evaluating the model’s performance multiple times to obtain a more accurate estimate of its accuracy.
4.2 Accuracy and Loss Analysis
The validation accuracies, training accuracies, and losses of each model are shown in Table 7. Our proposed model (MpoxSLDNet) exhibited a training accuracy of 97.53% and a training loss of 0.0862. The validation phase resulted in an accuracy of 94.56% and a loss of 0.1583 for the MpoxSLDNet model. Similarly, the training phase for DenseNet121 resulted in an accuracy of 85.86% and a loss of 0.3443. DenseNet121 achieved a validation accuracy of 84.38% with a validation loss of 0.3280. Moving on to ResNet50, the training accuracy and loss were recorded as 72.42% and 0.5333, respectively. The validation phase for ResNet50 resulted in an accuracy of 67.19% and a loss of 0.5602. Similarly, The training accuracy for VGG16 was recorded as 89.45%, with a corresponding training loss of 0.2404. VGG16 achieved a validation accuracy of 86.25% with a validation loss of 0.3080. In summary, the MpoxSLDNet models that we proposed demonstrated superior performance in terms of both training and validation accuracy, outperforming other models such as DenseNet121, ResNet50, and VGG16.
Table 6: Multiple Training accuracy and Testing accuracy for the MpoxSLDNet model.
@@ 4.2 Accuracy and Loss Analysis
Table 7: Evaluation of accuracy and loss for each model during training and validation.
@@
Table 8: Comparative Classification performance analysis of our proposed model vs. pre-trained model.
@@
4.3 Classification Report Analysis
The classification report gives a detailed breakdown of precision, recall, and F1 scores of each class. Using the classification report, we looked at precision, recall, and F1-score for each class to assess the proposed MpoxSLDNet model’s performance. Table 8 presents the classification report for each models used in this study, including DenseNet121, ResNet50, VGG16, and our proposed model (MpoxSLDNet). The classification report includes precision, recall, and F1-score values for each class, namely Monkeypox (0) and Non-Monkeypox (1), indicating the model’s performance in detecting monkeypox lesions. For DenseNet121, the precision for Monkeypox is 0.88, indicating that 88% of the samples classified as Monkeypox were correctly identified. The recall for Monkeypox is 0.80, indicating that 80% of the actual Monkeypox samples were correctly identified. The F1-score for Monkeypox is 0.84, providing a balanced measure of precision and recall. Similarly, for Non-Monkeypox class, the precision, recall, and F1-score values are provided. The same evaluation metrics are provided for ResNet50 and VGG16 models, highlighting their performance in detecting Monkeypox and Non-Monkeypox lesions. Our proposed model, MpoxSLDNet, demonstrates superior performance in both classes, with high precision, recall, and F1-score values of 0.94 for both the Monkeypox and Non-Monkeypox classes. This indicates that the MpoxSLDNet model achieved a high level of accuracy in correctly identifying both types of lesions. This classification report table allows for a comprehensive comparison of the models’ performance in detecting monkeypox lesions, providing insights into the precision, recall, and F1-score values for each class. These metrics help to evaluate and select the most effective model for accurate monkeypox lesion detection.
@@ 4.3 Classification Report Analysis
4.4 Confusion Matrix and AUC ROC Curve Analysis
@@
Confusion matrices are good tools for evaluating classification models. This confusion matrix shows how many correctly classified and misclassified samples there are for each class. Figure 15 shows the confusion matrix of our proposed MpoxSLDNet model. Here, 320 of the test images are evaluated. True Positive tells us how often the model correctly categorizes monkeypox. In the same way, True Negative denotes the model’s ability to correctly classify non-monkeypox as non-monkeypox. From Figure 15, we can observe that the MpoxSLDNet model accurately identifies 152 images as monkeypox (True positive) and 149 images as non-monkeypox (True Negative). According to Figure 15, we can conclude that our model recognized 301 images correctly out of 320. On the other hand, false positive means that the model incorrectly classifies non-monkeypox as monkeypox. Conversely, False negative shows the instances where the model mistakenly classifies monkeypox as non-monkeypox. We can also observe that 9 images of Non-Monkeypox were incorrectly classified as monkeypox positive(False Positive), and 10 images of Monkeypox were incorrectly classified as Non-Monkeypox(negative). It shows that our approach can accurately identify Monkeypox patients, with high true positives and true negatives. Additionally, we employed Receiver Operating Characteristic (ROC) curves to evaluate the classification performance which represents the true positive rate (TPR) plotted against the false positive rate (FPR), aiding in the detection of positive Monkeypox cases within the tested Monkeypox and Non-Monkeypox images. Here, AUC measures how much our model is capable of distinguishing between our two classes, Monkeypox and Non-Monkeypox. The AUC-ROC curve is displayed for all the employed models in Figure 16. According to the data provided in Table 9, the proposed model demonstrated an AUC-ROC score of 0.94. The AUC-ROC score of the DenseNet121, ResNet50, and VGG16 are 0.84, 0.68, and 0.86, respectively. It is observed that the best AUC-ROC score belongs to the MpoxSLDNet model. Hence, the MpoxSLDNet model is relatively effective at distinguishing between positive and negative instances.
@@
4.5 Overall Performance Evaluation and Comparative Analysis: Proposed Model vs. Pre-trained Models
@@
In this section, we present an overall comprehensive analysis and comparison of the performance between our proposed model, MpoxSLDNet, and pre-trained models. The evaluation is conducted using key metrics including accuracy, precision, recall, and F1-score, AUC score. As shown in Table 9, the highest precision score of 0.94 occupied by the MpoxSLDNet model indicates that the model is good at correctly identifying positive instances. On the other hand, as recall increases, monkeypox images are identified more frequently, and the opposite is also true. With a recall score of 0.94, the MpoxSLDNet model reflects its ability to correctly identify a significant number of monkeypox images. Apart from this, the F1 score combines precision and recall into one score. The F1-score achieves its highest value of 1, indicating excellent performance, and its lowest value of 0, indicating poor performance. The MpoxSLDNet model achieves the highest average F1 score of 0.94, indicating a superior balance between precision and recall compared to the pre-trained model. Furthermore, the accuracy metric measures how many instances are classified correctly. With an accuracy score of 0.94, the model was able to classify instances correctly. Figure 17 serves as a succinct overview of the model’s performance, allowing for easy comparison and identification of the top-performing model according to the selected metrics. The x-axis represents the metrics name as well as y-axis represents the corresponding values for precision, recall, F1-score, accuracy, and AUC score ranging from 0 to 1. The graph provides a visual representation of how each model performed in terms of these metrics. Higher values indicate better performance. Undoubtedly, MpoxSLDNet consistently demonstrates superior precision, recall, F1-score, and accuracy compared to the other models. This suggests that the MpoxSLDNet model surpasses the other models in accurately detecting monkeypox lesions.
@@
4.6 Evaluating the Performance of the Proposed Model in Comparison to Existing Research
Table 10 presents a comparative analysis between the existing study and the proposed system, both utilizing the Monkeypox Skin Lesion Dataset (MSLD). The accuracy scores, represented as percentages, reflect the models’ ability to accurately classify monkeypox lesions from non-monkeypox lesions. Our proposed MpoxSLDNet model demonstrates an accuracy score of 94.56%, surpassing all other models in accuracy. In comparison, the previously employed models exhibited accuracy scores ranging from 71.86% to 91.11%. Apart from, in this study, other pre-trained models like DenseNet121, ResNet50, and VGG16 demonstrate accuracy values ranging from 67.19% to 86.25%. This also highlights the superior accuracy achieved by MpoxSLDNet compared to the other models, showcasing its effectiveness in accurately classifying monkeypox lesions. Table 10 effectively summarizes the accuracy performance of the different models, highlighting the superior performance of MpoxSLDNet in accurately detecting and classifying monkeypox lesions.
Table 9: Overall performance of our proposed model vs. pre-trained model.

### v2

@@ III Methodology
III-A Data Collection
In this study, data are collected from a publicly accessible source known as the "Monkeypox Skin Lesion Dataset" [24]. The dataset includes two classes: Monkeypox and Non-Monkeypox, which are used for performing binary classification. The Monkeypox class contains 1428 skin images of monkeypox lesions. On the other hand, Non-Monkeypox class contains 1764 skin images of Non-Monkeypox lesions. The detailed data description is shown in Table II. Figure 3 illustrates a visual representation of the dataset’s image distribution, highlighting the proportion of monkeypox lesion images compared to non-monkeypox lesion images. This information is crucial for understanding the composition of the dataset and the relative representation of different lesion types, thus providing valuable insights into the data used for training and evaluation in our study. Figure 4 shows the Sample set of images of Monkeypox lesion and Non-Monkeypox lesion from the dataset.
 [^1] https://www.kaggle.com/datasets/nafin59/monkeypox-skin-lesion-dataset
TABLE II: Dataset Description.
@@ III-A Data Collection
Monkeypox In this dataset, the monkeypox class contains 1428 skin images of monkeypox lesions. Kaggle Datasets [24]
Non-Monkeypox Non-Monkeypox class contains 1764 Skin images of chickenpox lesions, measles lesions. Kaggle Datasets [24]
@@ III-B Data Pre-processing
This step is crucial for the deep learning model as it ensures that the input data is appropriately formatted for training the model, leading to enhanced model accuracy. In this step, collected data are passed through six pre-processing operations before it is fed into the model. These pre-processing operations include resizing images to uniform sizes, rescaling the pixel value, setting data_format, label encoding, in-place augmentation, and data [32]. All pre-processing operations are described here.
III-B1 Resizing Image
@@
Image resizing is the process of resampling or downsampling the image. Images in the dataset should have the same size as the CNN model. So, these stages are used to assure that all the images have the same size. This stage is also used to reduce the computational burden of training the model. This is accomplished with the aid of Python/TensorFlow.keras.preprocessing module. In this study, image dimensions are set to 224×224 dimension. The reason for selecting image size 224×224 is that this size allows better performance, more accurate training, and faster convergence of model [26].
III-B2 Rescale:
@@
Rescaling is the process of adjusting or normalizing the image data to a specific range of 0 to 1. ImageDatagenerator class is used to transform the pixel values of images from the original range of 0 to 255 into the desired range of 0 to 1, which is more suitable for neural network models. Different image has different pixel value. If rescale operation is not applied, the model only encounters certain images which leads to overfitting. This operation ensures that each image is given equal importance and weight during the model’s training process. This operation can be accomplished by specifying the rescale argument which involves multiplying each pixel by a specific ratio to achieve the desired scaling. In this scenario, the ratio used for rescaling the pixel values is 1/255, which is approximately equal to 0.0039 [22].
III-B3 Data_format:
@@
In this study, data_format is set to "channels_last" which means that the last dimensions of the input tensor are channels. In this "channels_last" format image is depicted as a 3D tensor that corresponds to (height, width, and channels). In deep learning frameworks like TensorFlow and Keras, this format is the default value in Keras. During training and evaluating the model, this format ensures compatibility, efficient computation, and data manipulation [27].
III-B4 Label Encoding:
@@
Label encoding is a process of transforming categorical labels into numerical representations, enabling deep-learning predictive models to comprehend them effectively. In this study, the dict and zip method is used to create a mapping between the two classes ‘Monkeypox’ and ‘Non_Monkeypox’ and their corresponding numerical values 0 and 1. By using the dict method, we have created a dictionary in which the keys are the numerical values and the values are corresponding categorical labels. The zip method then makes a list of rows by combining the numerical values with the categorical labels. With the resulting dictionary, the categorical labels can be transformed into numerical values, allowing the model to utilize these numerical representations for making predictions. [25].
III-B5 Data Augmentation (in-place):
@@
In this pre-processing operation, random image transformations performed by in-place data augmentation (on-the-fly) have been applied, which means that the dataset size does not increase. In this type of data augmentation, at the time of training, our network observes unique variations of our data during each epoch and captures different patterns and features with each iteration. To transform the image, ImageDatagenerator, a Keras image prepossessing library, is used. Image transformation operations available in the ImageDatagenerator function include brightness_range, zoom range, horizontal flip, etc. Table III depicts the augmentation technique used in this study. Figure 5 shows the in-place augmentation process [28]. The process of applying in-place data augmentation is [28]:
Step-1: At first, augmentation parameters including zoom_range, brightness_range, fill_mode, and horizontal_flip are defined.
Step-2: An instance of an ImageDataGenerator class, specifying the augmentation parameters is also created.
Step-3: An original batch of images is passed to the ImageDataGenerator class.
Step-4: Then, the batches of augmented data are generated and returned to the calling method.
Step-5: After that, the model is trained on batches of images.
@@
TABLE III: Augmentation technique used in this study.
@@
III-B6 Data Visualization:
This stage requires sampling and reviewing the input data to observe the randomly transformed images and to verify that the transformations are being applied as expected. This stage is also used to confirm the readability of the input images. ImageDataGenerator library is used to generate a batch of randomly transformed images to visualize the transformed images. These transformed images are then visualized using the visualization library Matplotlib via 2D representation with the new picture dimensions [25]. Figure 6 shows the visualization of the transformed image sample of our dataset.
III-C Splitting Dataset
@@ III-D Model Building
TABLE IV: Summary of models used in this study.
@@
III-D1 Architecture of our Proposed MpoxSLDNet model :
@@
Figure 12 shows the graphical representation of fully connected layer. Then the last layer is 2 units dense layer with a sigmoid activation function producing a binary classification output. Table IV shows the summary of hyperparameters used in the proposed MpoxSLDNet model.
TABLE V: Hyper parameters for the proposed MpoxSLDNet model.
@@ IV Result Analysis
IV-A Assessing the MpoxSLDNet model’s performance through multiple runs
Table VI shows the training and validation accuracy obtained from running the proposed model three times for 20 epochs. The purpose of multiple runs is to mitigate the effects of randomness during the training phase and provide a more reliable evaluation of how well the model performs. Based on the Table VI, the MpoxSLDNet model attained an average training accuracy of 96.89% and an average validation accuracy of 93.08%. While the performance of the model is consistent across the three runs, there is some variation in the accuracy values. This highlights the importance of evaluating the model’s performance multiple times to obtain a more accurate estimate of its accuracy.
IV-B Accuracy and Loss Analysis
The validation accuracies, training accuracies, and losses of each model are shown in Table VII. Our proposed model (MpoxSLDNet) exhibited a training accuracy of 97.53% and a training loss of 0.0862. The validation phase resulted in an accuracy of 94.56% and a loss of 0.1583 for the MpoxSLDNet model. Similarly, the training phase for DenseNet121 resulted in an accuracy of 85.86% and a loss of 0.3443. DenseNet121 achieved a validation accuracy of 84.38% with a validation loss of 0.3280. Moving on to ResNet50, the training accuracy and loss were recorded as 72.42% and 0.5333, respectively. The validation phase for ResNet50 resulted in an accuracy of 67.19% and a loss of 0.5602. Similarly, The training accuracy for VGG16 was recorded as 89.45%, with a corresponding training loss of 0.2404. VGG16 achieved a validation accuracy of 86.25% with a validation loss of 0.3080. In summary, the MpoxSLDNet models that we proposed demonstrated superior performance in terms of both training and validation accuracy, outperforming other models such as DenseNet121, ResNet50, and VGG16.
TABLE VI: Multiple Training accuracy and Testing accuracy for the MpoxSLDNet model.
@@ IV-B Accuracy and Loss Analysis
TABLE VII: Evaluation of accuracy and loss for each model during training and validation.
@@
TABLE VIII: Comparative Classification performance analysis of our proposed model vs. pre-trained model.
@@
IV-C Classification Report Analysis
The classification report gives a detailed breakdown of precision, recall, and F1 scores of each class. Using the classification report, we looked at precision, recall, and F1-score for each class to assess the proposed MpoxSLDNet model’s performance. Table VIII presents the classification report for each models used in this study, including DenseNet121, ResNet50, VGG16, and our proposed model (MpoxSLDNet). The classification report includes precision, recall, and F1-score values for each class, namely Monkeypox (0) and Non-Monkeypox (1), indicating the model’s performance in detecting monkeypox lesions. For DenseNet121, the precision for Monkeypox is 0.88, indicating that 88% of the samples classified as Monkeypox were correctly identified. The recall for Monkeypox is 0.80, indicating that 80% of the actual Monkeypox samples were correctly identified. The F1-score for Monkeypox is 0.84, providing a balanced measure of precision and recall. Similarly, for Non-Monkeypox class, the precision, recall, and F1-score values are provided. The same evaluation metrics are provided for ResNet50 and VGG16 models, highlighting their performance in detecting Monkeypox and Non-Monkeypox lesions. Our proposed model, MpoxSLDNet, demonstrates superior performance in both classes, with high precision, recall, and F1-score values of 0.94 for both the Monkeypox and Non-Monkeypox classes. This indicates that the MpoxSLDNet model achieved a high level of accuracy in correctly identifying both types of lesions. This classification report table allows for a comprehensive comparison of the models’ performance in detecting monkeypox lesions, providing insights into the precision, recall, and F1-score values for each class. These metrics help to evaluate and select the most effective model for accurate monkeypox lesion detection.
@@ IV-C Classification Report Analysis
IV-D Confusion Matrix and AUC ROC Curve Analysis
@@
Confusion matrices are good tools for evaluating classification models. This confusion matrix shows how many correctly classified and misclassified samples there are for each class. Figure 15 shows the confusion matrix of our proposed MpoxSLDNet model. Here, 320 of the test images are evaluated. True Positive tells us how often the model correctly categorizes monkeypox. In the same way, True Negative denotes the model’s ability to correctly classify non-monkeypox as non-monkeypox. From Figure 15, we can observe that the MpoxSLDNet model accurately identifies 152 images as monkeypox (True positive) and 149 images as non-monkeypox (True Negative). According to Figure 15, we can conclude that our model recognized 301 images correctly out of 320. On the other hand, false positive means that the model incorrectly classifies non-monkeypox as monkeypox. Conversely, False negative shows the instances where the model mistakenly classifies monkeypox as non-monkeypox. We can also observe that 9 images of Non-Monkeypox were incorrectly classified as monkeypox positive(False Positive), and 10 images of Monkeypox were incorrectly classified as Non-Monkeypox(negative). It shows that our approach can accurately identify Monkeypox patients, with high true positives and true negatives. Additionally, we employed Receiver Operating Characteristic (ROC) curves to evaluate the classification performance which represents the true positive rate (TPR) plotted against the false positive rate (FPR), aiding in the detection of positive Monkeypox cases within the tested Monkeypox and Non-Monkeypox images. Here, AUC measures how much our model is capable of distinguishing between our two classes, Monkeypox and Non-Monkeypox. The AUC-ROC curve is displayed for all the employed models in Figure 16. According to the data provided in Table IX, the proposed model demonstrated an AUC-ROC score of 0.94. The AUC-ROC score of the DenseNet121, ResNet50, and VGG16 are 0.84, 0.68, and 0.86, respectively. It is observed that the best AUC-ROC score belongs to the MpoxSLDNet model. Hence, the MpoxSLDNet model is relatively effective at distinguishing between positive and negative instances.
@@
IV-E Overall Performance Evaluation and Comparative Analysis: Proposed Model vs. Pre-trained Models
@@
In this section, we present an overall comprehensive analysis and comparison of the performance between our proposed model, MpoxSLDNet, and pre-trained models. The evaluation is conducted using key metrics including accuracy, precision, recall, and F1-score, AUC score. As shown in Table IX, the highest precision score of 0.94 occupied by the MpoxSLDNet model indicates that the model is good at correctly identifying positive instances. On the other hand, as recall increases, monkeypox images are identified more frequently, and the opposite is also true. With a recall score of 0.94, the MpoxSLDNet model reflects its ability to correctly identify a significant number of monkeypox images. Apart from this, the F1 score combines precision and recall into one score. The F1-score achieves its highest value of 1, indicating excellent performance, and its lowest value of 0, indicating poor performance. The MpoxSLDNet model achieves the highest average F1 score of 0.94, indicating a superior balance between precision and recall compared to the pre-trained model. Furthermore, the accuracy metric measures how many instances are classified correctly. With an accuracy score of 0.94, the model was able to classify instances correctly. Figure 17 serves as a succinct overview of the model’s performance, allowing for easy comparison and identification of the top-performing model according to the selected metrics. The x-axis represents the metrics name as well as y-axis represents the corresponding values for precision, recall, F1-score, accuracy, and AUC score ranging from 0 to 1. The graph provides a visual representation of how each model performed in terms of these metrics. Higher values indicate better performance. Undoubtedly, MpoxSLDNet consistently demonstrates superior precision, recall, F1-score, and accuracy compared to the other models. This suggests that the MpoxSLDNet model surpasses the other models in accurately detecting monkeypox lesions.
@@
IV-F Evaluating the Performance of the Proposed Model in Comparison to Existing Research
Table X presents a comparative analysis between the existing study and the proposed system, both utilizing the Monkeypox Skin Lesion Dataset (MSLD). The accuracy scores, represented as percentages, reflect the models’ ability to accurately classify monkeypox lesions from non-monkeypox lesions. Our proposed MpoxSLDNet model demonstrates an accuracy score of 94.56%, surpassing all other models in accuracy. In comparison, the previously employed models exhibited accuracy scores ranging from 71.86% to 91.11%. Apart from, in this study, other pre-trained models like DenseNet121, ResNet50, and VGG16 demonstrate accuracy values ranging from 67.19% to 86.25%. This also highlights the superior accuracy achieved by MpoxSLDNet compared to the other models, showcasing its effectiveness in accurately classifying monkeypox lesions. Table X effectively summarizes the accuracy performance of the different models, highlighting the superior performance of MpoxSLDNet in accurately detecting and classifying monkeypox lesions.
TABLE IX: Overall performance of our proposed model vs. pre-trained model.